in an examination consisting of 15 questions,8 marks are awarded for a correct answer and 4 marks are deducted for a wrong answer,how many correct answer are needed to score 72 marks in this examination?
Answer by mananth(16946)

correct answer-----------x
wrong answer-------------y
x+y=15
8x-4y=72
1.00 x + 1.00 y = 15.00 .............1
Total value
8.00 x -4.00 y = 72.00 .............2
Eliminate y
multiply (1)by 4.00
Multiply (2) by 1.00
4.00 x 4.00 y = 60.00
8.00 x -4.00 y = 72.00
Add the two equations
12.00 x = 132.00
/ 12.00
x = 11.00
plug value of x in (1)
1.00 x + 1.00 y = 15.00
11.00 + y = 15.00
y = 15.00 -11.00
y = 4.00
y = 4.00
11 correct answers